HIGH-RESOLUTION LASER-INDUCED FLUORESCENCE (LIF) SPECTROSCOPY OF THE DEUTER-ATED ISOTOPOMERS OF THE METHOXY RADICAL

\underline{JINJUN LIU}, JOHN T. YI and TERRY A. MILLER, Laser Spectroscopy Facility, Department of Chemistry, The Ohio State University, 120 W. 18th Avenue, Columbus, Ohio 43210.

Our high-resolution laser-induced fluorescence (LIF) spectroscopy apparatus has been improved to study the deuterated isotopomers of the methoxy radical under jet-cooled conditions (T∼2K). Doppler-free absorption spectra of iodine molecule were recorded simultaneously for absolute calibration,^{a} while relative calibration was done using the fringes of a new etalon. The spectral resolution (FWHM) for methoxy was limited to ∼300MHz by residual Doppler broadening and unresolved hyperfine structure. The center of spectral lines could be determined to ∼1/5 the FWHM with a comparable absolute accuracy (1σ∼50MHz). The 3_{0}^{2} and (6′)_{0}^{1} bands of the Ã^{2}A_{1}–X̃^{2}E_{3/2} transition for CHD_{2}O and CH_{2}DO have been recorded, as well as the 3_{0}^{2} and 6_{0}^{1} bands for CH_{3}O and CD_{3}O. Rotational and fine structure for all bands has been resolved and assigned. Details of the analyses will be presented in the following talks.

aH. Kato, Doppler-Free High Resolution Spectral Atlas of Iodine Molecule, Japan Society for the Promotion of Science, (2002).